Plese help:
Ricardo is 4 years less than twice as old as his sister. The sum of their ages is 20.How old is ricardo?
4-2x=20
-4 -4
________
2x = 16
2 2
x=8
This is what I got. Is it correct?

Answer by ptaylor(2198)

WHAT IS x?????
YOU SURE ARE ON THE RIGHT TRACK!!! I HAVE FOUND THAT LAYING OUT THESE WORD PROBLEMS AS I HAVE DONE BELOW MAKES THEM A LOT EASIER TO SOLVE BECAUSE YOU CAN ALWAYS SEE WHERE YOU ARE AND WHERE YOU ARE GOING.

Let x=Ricardo's Sister's age
Then 2x-4=Ricardo's age (we are told this)
Now we are also told that the sum of their ages is 20. So:
x+2x-4=20 add 4 to both sides
x+2x-4+4=20+4 collect like terms
3x=24 divide both sides by 3
x=8---------------------------------Ricardo's sister's age
2x-4=2*8-4=12 ------------------------Ricardo's age

CK
12+8=20
20=20
Also: 12 is 4 less than 2 times 8
Hope this helps---ptaylor